Meeting Information

  • 1930 UT (1330 MDT)
  • AOC 317; IP:; phone: +1 575-835-7370



  • Issues from the weekend; break down of problems and any lessons learned.
  • Report from the User's Committee meeting (Claire).
  • Report from GB Conference (Bryan)
  • Staff Travel:
    • Joe: 19 - 24 May; back in office tomorrow
    • Emmanuel: 21 May - 7 June
    • Rick: 25-29 May
    • Joe L: 27 May -3 June
    • Michael: 27 May - 1 June
    • Gustaaf: 1 - 4 June; 18 June; 21 June
    • Eric: 4 - 19 June
    • James, Jacqueline: leave this weekend!

Commissioning review/reports

Notes for week of 10 May 2010 (Black: old but still relevant; Blue: new)

  • Meta summary
    • Multi-config SBs in regular use; note 'overhead' in configuration changes; Must add 30 seconds to each configuration change - Gustaaf's group will check projects for this until relaxed, in particular, reference pointing must expand to 3 minutes..

Mode OPT Scheduling Executor/Corr Calibration strategy Archive CASA AIPS Data Distribution Other Overall Status
  wide band modes available only in webtest OST (reconciling status database)

OPT selected limits not possible
Running 20 Hz
18-May CM version

21-May CBE version

    - needs check on short pol'n runs Need decision on large (i.e., RSRO) data set distribution media   Operational for OSRO1/OSRO2
30s overhead with OSRO configuration changes
RSRO beginning

Critical Commissioning

System Verification

  • Array/Receivers update/plans(Rick)
  • WIDAR update/plans (Michael)
    • Wideband status
    • Issues: MCAF; disk space; fallout of premature script exits; TelCal & multicast
  • Quick Software Systems review:
    • OPT (Lorant):webtest moving to production
    • Executor (Barry):
    • Scheduling (Joan, Bryan):
    • Correlator (Michael):delays; ea24 oddities
    • CASA (Juergen/Nick):
      • Developer Meeting briefing
    • AIPS (Eric):
  • Hardware notes:
    • Wideband checkouts: Dhawan, Krauss, Owen, Perley, Brogan, Hunter, Chomiuk, Marvil
      • Filler issues
      • Amount of time lost at reconfiguration(s)
      • Zeroes apart from reconfigurations
      • Data labels
      • RFI and tuning issues
      • Gain stability
      • Subbands 0 and 7
      • Images and other results

Band Checkouts (ECDATA-55) -- No updates.

Encompasses principally core operation at those frequencies followed by detailed calibration strategies, flagging.

  • P Band Status: %Red%Testing
    • Antennas with no P-band Rx: 7, 16, 22, 24 (no feed); 1, 4, 5 Rxs are not working.
  • L Band Status: Ready
    • Frequency labeling (ECDATA-9, ECDATA-10): Ok
    • Sky orientation (ECDATA-10): Ok
    • Gain stability (ECDATA-23): Ok; some bandpass phase evolution seen; follow-up on really blank sky.
    • Tuning: updated.
    • Antennas with no L-band: EA06, EA20, EA26
    • Integrating down (ECDATA-23):
      • Averaging in time is ok; averaging between subbands has problems.
      • L band (really) blank field (ECDATA-40)
    • L band, HI checkout (ECDATA-47)

  • S Band Status: Ready.
    • S band edge tests (similar result to C for tunings near 4 GHz; potentially some satellite downlink RFI; review of aliasing/out-of-band intermodulation is under review (Rick)
    • Frequency labeling (ECDATA-24): Ok
    • Sky orientation (ECDATA-18): Ok
    • Gain stability (ECDATA-24): Ok (no jumps seen)
    • Tuning (ECDATA-24): Reviewing default frequency based on birdies; recommendation to shift to multiple of 128 MHz .
    • Integrating down (ECDATA-18): Ok
    • Antennas with S band: EA08, EA09, EA15, EA22, EA24, EA27, EA28

  • C Status: Ready
    • Frequency labeling: Ok. Will get more information from Eric's analysis.
    • Sky orientation (ECDATA-35): Ok
    • Gain stability (ECDATA-35): Ok
    • Tuning: Low frequency tuning issue; stepped up for last AH1004.
    • Integrating down (ECDATA-25):
    • Antennas with no C-band: EA13

  • X Band Status: Ready
    • Frequency labeling : ECDATA-17): Ok
    • Sky orientation (ECDATA-17): Ok
    • Gain stability (ECDATA-26): Ok (no jumps)
    • Tuning: No issues
    • Integrating down (ECDATA-26): Ok
    • Antennas
    • RFI: none seen

  • K Band Status: Ready

  • Q Band Status: Ready.
    • Need Q band suite of tests.
    • Antennas with no Q-band: EA10, EA12, EA17

Observing Mode testing

  • Continuum - ok
  • Spectral Line: (Ott); ECSV-19
    • Bandpass stability: ECSV-35 (on hold due to multiple configuration changes).
    • Bandpass tests: ECSV-92
  • Polarimetry Status: (Myers); ECSV-29
    • C (ECDATA-12): Ready
    • L (ECDATA-42): Ready
    • Combined Low-High Pol checkout test (TPOL0003_sb1378180): Ready for obs (weekend source); pending 30 s overhead issue.
    • Eric reviewing AH1004 data for AIPS
    • Rick: Pol'n tests (receiver rotation) to be scheduled
  • Planetary Bryan away; need planning.
  • Advanced
    • Wideband L (AB1345 - ) Sanjay
    • Solar (Rick) status and path forward. Working with L, Ku (special Rx setups in these two bands). TP tests coming.
    • Phased Array:
      • Phased array for pulsars
    • Subarray handling
      • Single-dish VLBI
    • Astrometry
    • OTF Mosaicing

Observing support (e2e capability)

  • OSRO Testing (Gustaaf); ECSV-18 - needs to be pushed into JIRA.
  • Flux Calibrator Models (Amy); ECSV-43
    • Rupen proposal

Advanced Science Support

OSRO Proposal Testing (ECDATA-34)


  • Observations
    • RSRO Readiness (Cycle1) Master List
    • Further 1 GHz tests
    • Hoping for 2 GHz/16 subbands this week
    • How to schedule your program
      • "Priming the pump" with OSRO
      • Scan lengths, and duration vs. dwell
      • Time lost at configuration changes
      • Referenced pointing
      • Wise choice of frequencies & subbands
      • Production vs. webtest
      • Communicating additional information & correlator setups
  • Residents
    • Manuel Aravena (POC: Chris Carilli); area(s): High frequency checkout (reference pointing tests); OSRO data validation
      • This week:
    • Crystal Brogan (POC: Claire Chandler); area(s): High frequency observing; demo science
      • This week:
    • Laura Chomiuk (POC: Miller Goss; area(s): Correlator checkout; Spectral line commissioning; OSRO data validation
      • This week:
    • Adam Deller (POC: Michael Rupen); area(s): phased array system design and implementation
      • This week:
    • Todd Hunter (POC: Claire Chandler; area(S): science support software interfaces for users and high frequency observing; band/polarization checkout.
      • This week:
    • Joe Lazio (POC: Claire Chandler): Low frequency observing.
      • This week:
    • James Miller-Jones (POC: Claire Chandler; area(s): demo science use cases
      • Leaving Saturday
    • Jacqueline van Gorkom (POC: Miller Goss; area(s): Correlator checkout; spectral line commissioning)
      • Leaving Saturday
    • ---
    • Dave Green (POC: Sanjay Bhatnagar): area(s): algorithm tests for wideband imaging (currently away).


  • [Rupen]: Develop daily/post-maintenance day checklists for the system. To be completed as part of stress tests.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley]: Develop set of frequency accuracy tests (X band is complete).
  • [Perley]: Perform a frequency sweep test at different bands to provide to engineers for review (Q2).
  • [Wrobel/Butler]: Make use of API criteria consistent throughout tools (OPT->OST). List of actions organized. In process.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Subband 0 problems: Being reviewed on Wednesday morning at the WIDAR meeting. Butler to report.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: Occasional zero-amplitude visibilities (needs software fix - what level). UVLOD and FITLD can handle this in AIPS; flagdata clip is effective in CASA; still need to understand and fix at the source.
      • Low Priority [TBD]: Occasional pure-noise scans.
      • Low Priority [Rupen/Pokorny]: Delay clunking; software fix pending.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.


-- JosephMcMullin - 24 May 2010
Topic revision: r2 - 2010-05-24, mrupen
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback